(y^3+216)/(y^2-36)*(y-6)/6y
1
Expert's answer
2010-06-30T05:12:56-0400
(y^3+216)/(y^2-36)*(y-6)/6y = (y + 6)(y^2 - 6y + 36)/(y - 6)(y+6) * (y-6)/6y = (y^2 - 6y + 36)/(y - 6) * (y-6)/6y = (y^2 - 6y + 36)/6y = y/6 - 1 + 6/y